# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v2

@@ IV.4 Robustness
We investigate the robustness of the double dressed entangling gate against symmetric (asymmetric) detunings $\delta_{1}$ ($\delta_{2}$) of $\Omega^{\text{Amp}}_{1}$ w.r.t. $\nu-\epsilon$ by measuring the entangled state negativity at a gate time of 313 $\mathrm{\mu}$s with $\Omega_{1}^{\text{Phase}}=2\pi\times 94.8$ kHz. Fig. 6(a) shows the effect of a mismatch $\delta_{1}$ between the physical Rabi frequency $\Omega_{1}^{\text{Amp}}$ for both ions and the phase modulation frequency $\Omega_{1}^{\text{Phase}}$ ($\delta_{1}=\Omega_{1}^{\text{Phase}}-\Omega_{1}^{\text{Amp}}$). Fig. 6(b) shows the impact of a mismatch of the physical Rabi frequencies between the two ions ($\delta_{2}=\Omega_{1}^{\text{Amp},1}-\Omega_{1}^{\text{Amp},2}$), while $\Omega_{1}^{\text{Amp},1}=\Omega_{1}^{\text{Phase}}$ holds.
@@
For $\delta_{1}\approx 3\leavevmode\nobreak\ \%$ and $\delta_{2}\approx 8\leavevmode\nobreak\ \%$ of $\Omega_{1}^{\text{Phase}}$, the measured negativity stays above $80\leavevmode\nobreak\ \%$ of the optimal negativity, showing the robustness of the entangling gate against such mismatches of the RF field amplitude. Similar results were observed for other RF driven MS-type gates [60], quantified in terms of fidelity. For small detunings, as typically occur in experiments, a quadratic dependence between negativity and detuning is observed, showing the stability against small detunings. The robustness of the gate, made it possible to scan a wide range of $\delta_{1,2}$, exceeding typical uncertainties during the experiments by two orders of magnitude.
@@
Figure 6: Negativity of the two qubit state at the gate time depending on the symmetric (asymmetric) detuning $\delta_{1}$ ($\delta_{2}$) in the dressed state basis indicating the gate’s robustness. (a) $\delta_{1}=\Omega_{1}^{\text{Phase}}-\Omega_{1}^{\text{Amp}}$ symmetric for both ions. (b) $\delta_{2}=\Omega_{1}^{\text{Amp},1}-\Omega_{1}^{\text{Amp},2}$while $\Omega_{1}^{\text{Amp},1}=\Omega_{1}^{\text{Phase}}=2\pi\times$94.8 kHz.
@@ V Conclusion
V.1 Phase modulation of RF drive
For implementing the two-qubit entangling gate described in this work, a single, phase-modulated RF field is applied to each qubit. In addition, we study the effect of this phase-modulated driving field on single qubit coherence (section III). Alternatively, for these purposes an amplitude-modulated field could be used [44, 57]. We discuss the advantages of phase modulation as compared to amplitude modulation for these purposes in what follows.
Cohen et al. [57] put forth two schemes for the generation of double-dressed states: (a) applying a second RF driving field that is on resonance with the dressed-state energy splitting determined by the first field driving the bare qubit states, or (b) by applying a second driving field that is detuned from the bare-state energy splitting by exactly the Rabi frequency. Here, we focus exclusively on case (b) due to its straightforward implementation with a single RF signal generated by an AWG.
In case (b), two options are possible: b1) A phase-modulated implementation, which is used in this work, and b2) an amplitude-modulated implementation. In both of these approaches, the modulation of a single driving field is employed to generate sidebands around the carrier frequency instead of adding an independent driving field as in case (a). In this case the modulated signal, as guaranteed by the Jacobi-Anger expansion (e.g. [61]), ensures that the frequency components utilized to construct the gate are of equal amplitude relative to that of the carrier signal.
For phase modulation, the modulation depth $\Omega_{2}/\Omega_{1}$ can be as precise as the phase resolution of the signal applied to the qubit in the $\nu\approx 2\pi\times 12.6\leavevmode\nobreak$GHz regime. In this regime, a phase resolution of less than $\triangle\varphi=3.5\times 10^{-2}\leavevmode\nobreak$rad can be achieved through experimental means. Accordingly, at the modulation frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, a phase resolution of $\triangle\varphi=\triangle\varphi\cdot\Omega_{1}/\nu=3.5\times 10^{-7}\leavevmode\nobreak$rad can be achieved. As a consequence of the aforementioned resolution, a phase-modulated driving field permits the amplitude of the sidebands to be precisely calibrated with a relative step size of $10^{-5}$. Furthermore, the phase-modulated approach is immune against the non-linearity of RF amplifiers used for signal processing (see Appendix C). For these reasons, we believe that the phase-modulated implementation is a more promising avenue for future developments.
V.2 Summary and outlook
The best entangling gate reported for RF-controlled ions [17], although high in fidelity, is still an order of magnitude slower (740 $\mathrm{\mathrm{\mu}s}$) than its optical counterparts.
In this work, we lay the theoretical and experimental framework for implementing RF driven double dressed state entangling gates with further increasing gate speed. Using this novel gate scheme, we create the Bell states $|\Phi^{+}\rangle$ and $|\Psi^{+}\rangle$ simply by choosing the appropriate interaction time of qubits with a phase-modulated driving field. We achieved fidelities up to $98^{+2}_{-3}\leavevmode\nobreak\ \%$ with gate time of $\leq 313\leavevmode\nobreak\ \mathrm{\mu}\text{s}$, using readily available experimental parameters. To our knowledge, this is the first experimental realization of an entangling gate based on double-dressed qubits.
The qubits are double dressed by a single phase-modulated RF field which acts as continuous dynamical decoupling driving the entangling operation, and at the same time, protects the qubits’ coherence. Thus, this gate is intrinsically robust against RF field amplitude fluctuations, as well as magnetic field fluctuations that otherwise would adversely affect magnetically sensitive hyperfine-qubits. Although the concrete implementation is platform-specific, double-dressing using phase modulation can be achieved in other physical systems.
@@ V.2 Summary and outlook
We note that RF driven Mølmer-Sørensen-type gates with a gate time comparable to the results achieved in this work have recently been reported in [63].
@@ Acknowledgments
D.C. acknowledges the support of the Clore Scholars programme and the Clore Israel Foundation. M.N, P.H., P.B., D.N. and C.W. acknowledge the support of the EU Horizon 2020 Project No.820314 (microQC) and the German Federal Ministry of Education and Research under grant number 13N15521 (MIQRO). M.N, P.H., and D.N. thank Matthias Kleinmann for fruitful discussions.
@@
Appendix A Double dressed state of a single Qubit
In the main text, we give a brief summary of the theoretical concepts used to interpret the experimental results. Here, in the appendix, we give a more complete account of these. In order to make this appendix self-contained for the convenience of the reader, we repeat here expressions that are also used in the main text.
We consider driving a single ion co-trapped with N-1 ions in a 1-D harmonic potential. The Hamiltonian that describes two hyperfine states, given that the ion is subjected to a linear magnetic field gradient, is [9]
@@ Appendix A Double dressed state of a single Qubit
where $\omega_{0}$ is the transition frequency between the two hyperfine states, $\nu$ is the motional mode frequency, $\eta$ is the effective Lamb-Dicke parameter, $\sigma_{i}$ is the Pauli matrix in the $i$-th direction ($i=x,y,z$) and $b/b^{\dagger}$ is the lowering/raising operator of the motional quanta.
@@
Here $\triangle z$ is the extension of the motional ground state wave function. The qubit’s addressing frequency, $\omega(z)$, depends on the external magnetic field as described by the Breit-Rabi formula [64] for ${}^{171}\text{Yb}^{+}$
@@
Here, $E_{\text{HFS}}$ is the hyperfine splitting between levels with total angular momentum $F=I+1/2$ and $F=I-1/2$ in zero magnetic field, $\mu_{B}$ is Bohr’s magneton, $m_{e}$ and $m_{p}$ indicate the electron and proton mass, respectively, $g_{J}$ and $g_{I}$ are the electronic and nuclear g-factor, and the magnetic field is given by $B(z)=B_{0}+\delta_{z}B$. In equation 14 the linear Zeeman effect is considered.
We drive the ion near-resonance with a phase modulated field such that
@@
where $\Omega_{1}$ is the Rabi frequency of the RF drive and $\Omega_{2}$ is a parameter quantifying the phase modulation amplitude.
The total Hamiltonian, $H+H_{D}$, in the rotating frame w.r.t. $H_{0}=\frac{\omega_{0}}{2}\sigma_{z}+\frac{\Omega_{2}}{2}\sigma_{z}\cos(\Omega_{1}t)$, can be written as
@@
where we made the RWA assuming $\Omega_{1}\ll 2\omega_{0}$ and $\Omega_{2}\ll\Omega_{1}$.
We define the dressed basis operators by the canonical transformation $S_{x}=-\sigma_{z},\ S_{y}=\sigma_{y},S_{z}=\sigma_{x}$. The Hamiltonian in Eq. (19) then takes the form,
@@
As in other continuous dynamical decoupling schemes, DC noise components perpendicular to the dressed energy gap are suppressed as long as they are much smaller than $\Omega_{1}$. Specifically, magnetic field noise that causes a shift of the bare energy gap $\omega_{0}\rightarrow\omega_{0}+\delta\omega_{0}$ is suppressed when $\delta\omega_{0}\ll\Omega_{1}$. Similarly, time dependent noise is suppressed, if the power spectral density of the noise at frequency $\Omega_{1}$ is small. We also note that the specific choice of phase modulation in Eq. (18) translates into an on-resonance drive on the dressed qubit in the rotating frame (Eq. (20)).
@@
where we made the RWA assuming $\Omega_{2}\ll 4\Omega_{1}$.
We define the double dressed states operators by the transformation $F_{z}=S_{x},\ F_{y}=S_{y},\ F_{x}=-S_{z}$. Consequently, the Hamiltonian in Eq. (21) can be written as
@@
The effective second drive in Eq. (20) translates into a second dressed energy gap in the double rotating frame (Eq. (22)). This suppresses noise sources perpendicular to the double dressed energy gap that survive in the rotating frame, Eq. (20). Specifically, amplitude fluctuations in the drive, $\Omega_{1}\rightarrow\Omega_{1}+\delta\Omega_{1}$ will cause the dressed energy gap to fluctuate, reducing the coherence time of the dressed qubit. These will be suppressed as long as $\Omega_{2}\gg\delta\Omega_{1}$.
Another noise source arises from the coupling of the internal qubit states to the motional states. To unfold this statement, we first take $\Omega_{2}=0$, which describes a Rabi experiment. Then Eq. (21) describes the undesired coupling to the motional states in the dressed basis. Assuming that $\Omega_{1}+\nu\gg\eta\nu$, the interaction can be written as
@@
The Hamiltonian in Eq. (23) creates an effective shift on the dressed basis that depends on the motional state of the ion. In the limit $\eta\nu\ll|\nu-\Omega_{1}|$ it can be described by [65]
@@
Therefore, if the motional state is far from the motional ground state we will observe a decay in the Rabi oscillations due to this state dependent frequency shift.
The effective second drive decouples this interaction as well when $\Omega_{2}\gg\eta\nu$ and as long as $|\Omega_{1}-\nu|\ll\Omega_{2}\ll\Omega_{1}+\nu$. In the laser-free trapped ion setting, using a static magnetic field gradient, this allows to counter further decoherence induced by the motional states.
@@ Appendix B Fidelity estimation for higher gradients
In the main text we claim that a higher magnetic field gradient will result in a faster gate with higher fidelity. In what follows, we estimate the different mechanisms that affect the fidelity in higher gradients. First, heating of the motional modes is considered. As was shown in the original work by Mølmer and Sørensen [48] this translates to an effective dephasing rate of
@@
Note, that the coupling of the internal ion states to the motional mode in Ref. [48] is of the form $\sigma_{y}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$ while for the double-dressed states we have $F_{z}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$. Therefore, in our case the dephasing is along the Z axis, compared to the Y axis for a MS gate.
Increasing the gradient will also increase the effect of the off-resonance unitary terms and the unitary infidelity accordingly. The issue is the quantum Stark shift that affects the double-dressed states, as it leaves the ions entangled to the phonons at the end of the gate. In the following we drive an estimate for this shift.
Assuming, that the Rabi frequency $\epsilon\ll\nu$, we can approximate Eq. (22) to
@@
Moving to the rotating frame w.r.t. the double dressed state energy $\frac{\Omega_{2}}{4}F_{z}$, (26) transforms to
@@
Note, that when $\Omega_{2}/2=\epsilon$, the flip-flop interaction is on-resonance. We suspect that this is the reason for the dip in the coherence times in Fig. 2. While the RWA $\epsilon\gg\eta\nu$ that justifies neglecting the rotating terms is not valid in the limit $\Omega_{2}\sim\epsilon=\eta\nu$, we still think it is helpful to understand the dynamics by the following argument. The dressed basis changes due to the Jaynes-Cummings interaction into a hybridized basis with the phonons. Then the expected coherence time can be approximated by the Jaynes-Cummings collapse time [67] $\sim 2\sqrt{2}/\eta\nu\sim 0.1\ \text{ms}$. Although deviations from this value are expected due to the crude nature of the approximation, we believe that the underlying reason for the dip in the coherence time is the same.
Returning to the operating regime of the gate, where $\left|\Omega_{2}/2\pm\epsilon\right|\gg\eta\nu/2$, the shift of the double-dressed state’s energy that depends on the phonon occupation can be estimated by the second-order effective Hamiltonian [65]
$H_{eff}$ $=\left(\frac{\eta\nu}{4}\right)^{2}\left[\frac{1}{\Omega_{2}/2-\epsilon}+\frac{1}{\Omega_{2}/2+\epsilon}\right]F_{z}b^{\dagger}b$ (29)
@@
where in the last transition we assumed that $\frac{\Omega_{2}}{2}\gg\epsilon=\nu-\Omega_{1}$. Denoting $\delta_{qss}=\frac{\left(\eta\nu\right)^{2}}{4\Omega_{2}}$, the gate dynamics can be understood by the following Hamiltonian
@@
Assuming we start at the motional ground state, the first order Magnus expansion suggests that we create two distinct coherent states depending on the ions’ state:
@@
while $\alpha_{01}=\alpha_{10}=0$. At the gate time we have
@@
$IF\approx$ $\frac{\left|\alpha_{00}(t_{g})\right|^{2}+\left|\alpha_{11}(t_{g})\right|^{2}}{4}=$ (33)
@@
$IF\approx\frac{\left(\delta_{qss}t_{g}\right)^{2}}{2}.$ (34)
For our current setup $\delta_{qss}\approx 0.03\leavevmode\nobreak\ \text{kHz}$, therefore, the expected infidelity is about $0.2\leavevmode\nobreak\ \%$. This estimation fits well with our simulations starting at the motional ground state. However, in our experiment we start with $\bar{n}\approx 0.6$ which results in a significant deviation from this value and the unitary infidelity reaches about $1.5\%$. Since this term increases approximately quadratically with the magnetic field gradient, it is expected to be the dominant source of error in the gate.
There are two ways to mitigate this error. The first is to work with the stretch mode instead of the COM mode. This will allow to increase $\Omega_{2}$ significantly (by a factor of $\sqrt{3}$ similar to the mode’s frequency $\nu_{\text{str}}=\sqrt{3}\nu_{\text{COM}}$) and therefore suppress this interaction efficiently. In particular, the Lamb-Dicke parameter (Eq. 14) decreases by $3^{-3/4}$, and the frequency $\delta_{qss}\sim(\eta\nu)^{2}/\Omega_{2}$ is reduced by $3^{-1}$. This scaling will lower the infidelity by $3^{-3/2}$ while the gate time only increases by a factor of $3^{1/4}$ due to $t_{g}\sim 1/(\eta\nu)$.
A second method for reducing this interaction is to employ additional pulse shaping techniques. Our simulation of the gate with the current experimental parameters and a standard Gaussian ramp of $\Omega_{1}$ demonstrates a notable reduction in unitary infidelity to approximately 0.1-0.2 %, while slightly increasing the gate duration to approximately 410-450 $\mathrm{\mathrm{\mu}s}$. It is expected that similar improvements can be achieved with higher gradients. More sophisticated pulse schemes are predicted to further enhance this reduction, potentially by an order of magnitude.

### v3

@@ IV.4 Robustness
We investigate the robustness of the double-dressed entangling gate against symmetric (asymmetric) detunings $\delta_{1}$ ($\delta_{2}$) of $\Omega^{\text{amp}}_{1}$ with respect to $\nu-\epsilon$ by measuring the negativity of the entangled state at a gate time of 313 µs with $\Omega_{1}^{\text{phase}}=2\pi\times 94.8$ kHz. Figure 6(a) shows the effect of a mismatch $\delta_{1}$ between the physical Rabi frequency $\Omega_{1}^{\text{amp}}$ for both ions and the phase-modulation frequency $\Omega_{1}^{\text{phase}}$ ($\delta_{1}=\Omega_{1}^{\text{phase}}-\Omega_{1}^{\text{amp}}$). Figure 6(b) shows the impact of a mismatch of the physical Rabi frequencies between the two ions ($\delta_{2}=\Omega_{1}^{\text{amp},1}-\Omega_{1}^{\text{amp},2}$), while $\Omega_{1}^{\text{amp},1}=\Omega_{1}^{\text{phase}}$ holds.
@@
For $\delta_{1}\approx 3\%$ and $\delta_{2}\approx 8\%$ of $\Omega_{1}^{\text{phase}}$, the measured negativity stays above $80\%$ of the optimal negativity, showing the robustness of the entangling gate against such mismatches of the rf field amplitude. Similar results were observed for other rf-driven MS-type gates [60], quantified in terms of fidelity. For small detunings, as typically occurs in experiments, a quadratic dependence between negativity and detuning is observed, showing stability against small detunings. The robustness of the gate makes it possible to scan a wide range of $\delta_{1,2}$, exceeding typical uncertainties during the experiments by 2 orders of magnitude.
@@
Figure 6: Negativity of the two-qubit state at the gate time depending on the symmetric (asymmetric) detuning $\delta_{1}$ ($\delta_{2}$) in the dressed-state basis, indicating the gate’s robustness. (a) $\delta_{1}=\Omega_{1}^{\text{phase}}-\Omega_{1}^{\text{amp}}$ symmetric for both ions. (b) $\delta_{2}=\Omega_{1}^{\text{amp},1}-\Omega_{1}^{\text{amp},2}$while $\Omega_{1}^{\text{amp},1}=\Omega_{1}^{\text{phase}}=2\pi\times$94.8 kHz.
IV.5 Comparison to double-drive schemes
For implementing the two-qubit entangling gate described in this work, a single phase-modulated rf field is applied to each qubit. In addition, we study the effect of this phase-modulated driving field on single-qubit coherence (Sec. III). Alternatively, for these purposes, an amplitude-modulated field could be used [44, 57]. We discuss the advantages of phase modulation as compared to amplitude modulation for these purposes in what follows.
Cohen et al. [57] put forth two schemes for the generation of double-dressed states: (a) applying a second rf driving field that is on resonance with the dressed-state energy splitting determined by the first field driving the bare qubit states, or (b) applying a second driving field that is detuned from the bare state energy splitting by exactly the Rabi frequency. Here, we focus exclusively on case (b) due to its straightforward implementation with a single rf signal generated by an AWG.
In case (b), two options are possible: (b1) a phase-modulated implementation, which is used in this work, and (b2) an amplitude-modulated implementation. In both of these approaches, the modulation of a single driving field is employed to generate sidebands around the carrier frequency instead of adding an independent driving field as in case (a). In this case, the modulated signal, as guaranteed by the Jacobi-Anger expansion (e.g., Ref. [61]), ensures that the frequency components utilized to construct the gate are of equal amplitude relative to that of the carrier signal.
For phase modulation, the modulation depth $\Omega_{2}/\Omega_{1}$ can be as precise as the phase resolution of the signal applied to the qubit in the $\nu\approx 2\pi\times 12.6\leavevmode\nobreak$GHz regime. In this regime, a phase resolution of less than $\triangle\varphi=3.5\times 10^{-2}\leavevmode\nobreak$rad can be achieved through experimental means. Consequently, at the modulation frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, a phase resolution of $\triangle\varphi=\triangle\varphi\cdot\Omega_{1}/\nu=3.5\times 10^{-7}\leavevmode\nobreak$rad can be achieved. As a consequence of the aforementioned resolution, a phase-modulated driving field permits the amplitude of the sidebands to be precisely calibrated with a relative step size of $10^{-5}$. Furthermore, the phase-modulated approach is immune to the nonlinearity of rf amplifiers used for signal processing (see Appendix C). For these reasons, we believe that the phase-modulated implementation is a more promising avenue for future developments.
@@ V Conclusion
The best entangling gate reported for rf-controlled ions [17], although high in fidelity, is still an order of magnitude slower (740 µs) than its optical counterparts.
In this work, we laid the theoretical and experimental framework for implementing rf-driven double-dressed-state entangling gates with a further increasing gate speed. Using this novel gate scheme, we created the Bell states $|\Phi^{+}\rangle$ and $|\Psi^{+}\rangle$ simply by choosing the appropriate interaction time of qubits with a phase-modulated driving field. We achieved fidelities up to $98^{+2}_{-3}\%$ with a gate time of less than or equal to 313µs, using readily available experimental parameters. To our knowledge, this work is the first experimental realization of an entangling gate based on double-dressed qubits.
The qubits were double dressed by a single phase-modulated rf field, which acted as continuous dynamical decoupling driving the entangling operation, at the same time protecting the qubits’ coherence. Thus, this gate was intrinsically robust against rf field amplitude fluctuations, as well as magnetic field fluctuations that otherwise would adversely affect magnetically sensitive hyperfine qubits. Although the concrete implementation is platform specific, double dressing using phase modulation can be achieved in other physical systems.
@@
We note that rf-driven Mølmer-Sørensen-type gates with a gate time comparable to the results achieved in this work have recently been reported in Ref. [63].
@@ Acknowledgments
D.C. acknowledges the support of the Clore Scholars program and the Clore Israel Foundation. M.N, P.H., P.B., D.N. and C.W. acknowledge the support of the EU Horizon 2020 Project No. 820314 (microQC) and the German Federal Ministry of Education and Research under Grant No. 13N15521 (MIQRO). M.N, P.H., and D.N. thank Matthias Kleinmann for fruitful discussions.
@@
Appendix A DOUBLE-DRESSED STATE OF A SINGLE QUBIT
In the main text, we give a brief summary of the theoretical concepts used to interpret the experimental results. Here, we give a more complete account of these concepts. In order to make this appendix self-contained for the convenience of the reader, we repeat expressions that are also used in the main text.
We consider driving a single ion cotrapped with N-1 ions in a 1D harmonic potential. The Hamiltonian that describes two hyperfine states, given that the ion is subjected to a linear magnetic field gradient, is [9]
@@ Appendix A DOUBLE-DRESSED STATE OF A SINGLE QUBIT
where $\omega_{0}$ is the transition frequency between the two hyperfine states, $\nu$ is the motional mode frequency, $\eta$ is the effective Lamb-Dicke parameter, $\sigma_{i}$ is the Pauli matrix in the $i$th direction ($i=x,y,z$), and $b$ and $b^{\dagger}$ are the lowering and raising operators, respectively, of the motional quanta.
@@
Here $\triangle z$ is the extension of the motional ground-state wave function. The qubit’s addressing frequency $\omega(z)$ depends on the external magnetic field as described by the Breit-Rabi formula [64] for ${}^{171}\text{Yb}^{+}$,
@@
Here, $E_{\text{HFS}}$ is the hyperfine splitting between levels with total angular momentum $F=I+1/2$ and $F=I-1/2$ in zero magnetic field; $\mu_{B}$ is Bohr’s magneton; $m_{e}$ and $m_{p}$ indicate the electron and proton mass, respectively; $g_{J}$ and $g_{I}$ are the electronic and nuclear $g$-factor; and the magnetic field is given by $B(z)=B_{0}+\delta_{z}B$. In Eq. (14), the linear Zeeman effect is considered.
We drive the ion near resonance with a phase-modulated field such that
@@
where $\Omega_{1}$ is the Rabi frequency of the rf drive and $\Omega_{2}$ is a parameter quantifying the phase-modulation amplitude.
The total Hamiltonian, $H+H_{D}$, in the rotating frame with respect to $H_{0}=\frac{\omega_{0}}{2}\sigma_{z}+\frac{\Omega_{2}}{2}\sigma_{z}\cos(\Omega_{1}t)$ can be written as
@@
where we chose the RWA assuming $\Omega_{1}\ll 2\omega_{0}$ and ${\Omega_{2}\ll\Omega_{1}}$.
We define the dressed-basis operators by the canonical transformation $S_{x}=-\sigma_{z},\ S_{y}=\sigma_{y},S_{z}=\sigma_{x}$. The Hamiltonian in Eq. (19) then takes the form
@@
As in other continuous dynamical decoupling schemes, dc noise components perpendicular to the dressed energy gap are suppressed as long as they are much smaller than $\Omega_{1}$. Specifically, magnetic field noise that causes a shift of the bare energy gap $\omega_{0}\rightarrow\omega_{0}+\delta\omega_{0}$ is suppressed when $\delta\omega_{0}\ll\Omega_{1}$. Similarly, time-dependent noise is suppressed if the power spectral density of the noise at frequency $\Omega_{1}$ is small. We also note that the specific choice of phase modulation in Eq. (18) translates into an on-resonance drive on the dressed qubit in the rotating frame [Eq. (20)].
@@
where we chose the RWA assuming $\Omega_{2}\ll 4\Omega_{1}$.
We define the double-dressed states operators by the transformation $F_{z}=S_{x},\ F_{y}=S_{y},\ F_{x}=-S_{z}$. Consequently, the Hamiltonian in Eq. (21) can be written as
@@
The effective second drive in Eq. (20) translates into a second dressed energy gap in the double rotating frame [Eq. (22)]. This drive suppresses noise sources perpendicular to the double-dressed energy gap that survive in the rotating frame, Eq. (20). Specifically, amplitude fluctuations in the drive, $\Omega_{1}\rightarrow\Omega_{1}+\delta\Omega_{1}$, will cause the dressed energy gap to fluctuate, reducing the coherence time of the dressed qubit. These fluctuations will be suppressed as long as $\Omega_{2}\gg\delta\Omega_{1}$.
Another noise source arises from the coupling of the internal qubit states to the motional states. To unfold this statement, we first take $\Omega_{2}=0$, which describes a Rabi experiment. Then, Eq. (21) describes the undesired coupling to the motional states in the dressed basis. Assuming that $\Omega_{1}+\nu\gg\eta\nu$, the interaction can be written as
@@
The Hamiltonian in Eq. (23) creates an effective shift on the dressed basis that depends on the motional state of the ion. In the limit $\eta\nu\ll|\nu-\Omega_{1}|$, it can be described by [65]
@@
Therefore, if the motional state is far from the motional ground state, we will observe a decay in the Rabi oscillations due to this state-dependent frequency shift.
The effective second drive decouples this interaction as well when $\Omega_{2}\gg\eta\nu$ and as long as $|\Omega_{1}-\nu|\ll\Omega_{2}\ll\Omega_{1}+\nu$. This, in the laser-free trapped-ion setting, using a static magnetic field gradient, we can further counter decoherence induced by the motional states.
@@ Appendix B FIDELITY ESTIMATION FOR HIGHER GRADIENTS
In the main text, we claim that a higher magnetic field gradient will result in a faster gate with higher fidelity. In what follows, we estimate the different mechanisms that affect the fidelity in higher gradients. First, heating of the motional modes is considered. As was shown in the original work by Mølmer and Sørensen [48], this translates to an effective dephasing rate of
@@
Note that the coupling of the internal ion states to the motional mode in Ref. [48] is of the form $\sigma_{y}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$, while for the double-dressed states, we have $F_{z}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$. Therefore, in our case, the dephasing is along the $z$ axis, compared to the $y$ axis for a MS gate.
Increasing the gradient will also increase the effect of the off-resonance unitary terms and the unitary infidelity accordingly. The issue is the quantum Stark shift that affects the double-dressed states, as it leaves the ions entangled to the phonons at the end of the gate. In the following, we derive an estimate for this shift.
Assuming that the Rabi frequency $\epsilon\ll\nu$, we can approximate Eq. (22) as
@@
Moving to the rotating frame with respect to the double-dressed state energy $\frac{\Omega_{2}}{4}F_{z}$, Eq. (26) transforms to
@@
Note that when $\Omega_{2}/2=\epsilon$, the flip-flop interaction is on resonance. We suspect that this is the reason for the dip in the coherence times in Fig. 2. While the RWA $\epsilon\gg\eta\nu$ that justifies neglecting the rotating terms is not valid in the limit $\Omega_{2}\sim\epsilon=\eta\nu$, we still think it is helpful to understand the dynamics by the following argument. The dressed basis changes due to the Jaynes-Cummings interaction into a hybridized basis with the phonons. Then the expected coherence time can be approximated by the Jaynes-Cummings collapse time [67], around $2\sqrt{2}/\eta\nu\sim 0.1\ \text{ms}$. Although deviations from this value are expected due to the crude nature of the approximation, we believe that the underlying reason for the dip in the coherence time is the same.
Returning to the operating regime of the gate, where $\left|\Omega_{2}/2\pm\epsilon\right|\gg\eta\nu/2$, the shift of the double-dressed-state energy that depends on the phonon occupation can be estimated by the second-order effective Hamiltonian, [65]
$H_{\text{eff}}$ $=\left(\frac{\eta\nu}{4}\right)^{2}\left[\frac{1}{\Omega_{2}/2-\epsilon}+\frac{1}{\Omega_{2}/2+\epsilon}\right]F_{z}b^{\dagger}b$ (29)
@@
where, in the last transition, we assumed that $\frac{\Omega_{2}}{2}\gg\epsilon=\nu-\Omega_{1}$. Denoting $\delta_{qss}=\frac{\left(\eta\nu\right)^{2}}{4\Omega_{2}}$, the gate dynamics can be understood by the following Hamiltonian:
@@
Assuming we start at the motional ground state, the first-order Magnus expansion suggests that we create two distinct coherent states depending on the ions’ state:
@@
while $\alpha_{01}=\alpha_{10}=0$. At the gate time, we have
@@
[…]
@@
$\text{IF}\approx\frac{\left(\delta_{qss}t_{g}\right)^{2}}{2}.$ (34)
For our current setup, $\delta_{qss}\approx 0.03\leavevmode\nobreak\ \text{kHz}$, therefore, the expected infidelity is about $0.2\%$. This estimation fits well with our simulations starting at the motional ground state. However, in our experiment, we start with $\bar{n}\approx 0.6$, which results in a significant deviation from this value, and the unitary infidelity reaches about $1.5\%$. Since this term increases approximately quadratically with the magnetic field gradient, it is expected to be the dominant source of error in the gate.
There are two ways to mitigate this error. The first is to work with the stretch mode instead of the c.m. mode. This approach will allow us to increase $\Omega_{2}$ significantly (by a factor of $\sqrt{3}$ because $\nu_{\text{str}}=\sqrt{3}\nu_{\text{c.m.}}$) and therefore suppress this interaction efficiently. In particular, the Lamb-Dicke parameter [Eq. 14] decreases by $3^{-3/4}$, and the frequency $\delta_{qss}\sim(\eta\nu)^{2}/\Omega_{2}$ is reduced by $3^{-1}$. This scaling will lower the infidelity by $3^{-3/2}$ while the gate time only increases by a factor of $3^{1/4}$ due to $t_{g}\sim 1/(\eta\nu)$.
A second method for reducing this interaction is to employ additional pulse-shaping techniques. Our simulation of the gate with the current experimental parameters and a standard Gaussian ramp of $\Omega_{1}$ demonstrates a notable reduction in unitary infidelity to approximately 0.1%-0.2%, while slightly increasing the gate duration to approximately 410-450 µs. It is expected that similar improvements can be achieved with higher gradients. More sophisticated pulse schemes are predicted to further enhance this reduction, potentially by an order of magnitude.